OpenStudy (anonymous):

2.What is the sum of the geometric sequence 4, 12, 36 … if there are 9 terms?

OpenStudy (anonymous):

@texaschic101 i tried to do this one on my own but i am lost

OpenStudy (texaschic101):

first you have to find the difference....do you know what the difference is ?

OpenStudy (anonymous):

yes it is 4

OpenStudy (texaschic101):

try again

OpenStudy (anonymous):

oh sorry it is 3

OpenStudy (texaschic101):

correct. now the sum formula is : sn = a1( 1 - r^n) / (1 - r) n = 9 (because we want total of 9 terms) r = 3 a1 = 4 (a1 is the first term) you understand all that so far ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

now lets sub our stuff into the formula... s9 = 4 ( 1 - 3^9) / (1 - 3) s9 = 4 (1 - 19683) / (-2) with me so far ? I had to use an exponent calculator

OpenStudy (anonymous):

yes i am following so far

OpenStudy (anonymous):

i am wrighting down each step that you are showing me so i can do it on my own

OpenStudy (texaschic101):

s9 = 4(- 19682) / (-2) s9 = - 78728 / -2 (remember when dividing by two negatives the answer will be positive) s9 = 39364 do you have any questions ?

OpenStudy (anonymous):

yes do i divide s9 into 39364?

OpenStudy (texaschic101):

no....s9 is just saying the sum of the 9 terms....your answer is 39364

OpenStudy (anonymous):

oh ok i understand

OpenStudy (texaschic101):

cool. You know I still can't remember this formula, I always have to look it up. I have so many formulas running in my head...lol

OpenStudy (anonymous):

yeah i know that feeling wjhy does there have to be so many formulas

OpenStudy (texaschic101):

I don't know, but there is way toooo many to remember

OpenStudy (anonymous):

so true

OpenStudy (anonymous):

thanks again:)

OpenStudy (texaschic101):

anytime :)